OpenStudy (kidthatbro8):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and The profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, …where x is the number of smart phone cases sold and y is the number of tablet cases sold. 3. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.

OpenStudy (igreen):

Oh, remember we changed this graph to \(y = -\dfrac{2}{3}x + 400\)?

OpenStudy (kidthatbro8):

yes.

OpenStudy (igreen):

Function notation is just changing y to f(x). Can you do that?

OpenStudy (igreen):

For example Function notation of \(y = x + 3\) is \(f(x) = x + 3\).

OpenStudy (kidthatbro8):

would it be f(x) = −2/3x + 400?

OpenStudy (kidthatbro8):

f(x) = -2/3x + 400, sorry, i don't know what went wrong up there.

OpenStudy (igreen):

Yep.

OpenStudy (igreen):

Open Study went wrong..lol. Just click on a different post and click back here..it'll go back to normal. Or you can reload the page.

OpenStudy (igreen):

The question already tells you what the function represents: "It represents the company's profit from cases last month."

OpenStudy (igreen):

That answers your question.

OpenStudy (kidthatbro8):

oh, thank you !

OpenStudy (igreen):

No problem.